40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 24 13:28:48 2014, Steve, Update, VAC, vent is completed ventiscompleted.pngatAtm.png
    Reply  Wed Sep 24 16:02:58 2014, Jenne, Update, VAC, vent is completed 
       Reply  Wed Sep 24 18:17:46 2014, ericq, Update, General, Alignment Restored 
       Reply  Thu Sep 25 08:21:15 2014, Steve, Update, VAC, vent day 1 ventDay1.png
          Reply  Mon Sep 29 09:11:25 2014, Steve, Update, VAC, vent day 5 RGA scan v77_bg-mag-d5.png
             Reply  Mon Sep 29 10:29:25 2014, Steve, Update, VAC, RGA is not running rgascanTimeStamp.png
Message ID: 10534     Entry time: Wed Sep 24 18:17:46 2014     In reply to: 10533
Author: ericq 
Type: Update 
Category: General 
Subject: Alignment Restored 

Interferometer alignment is restored

ASS has been run on each arm, recycling mirrors were aligned by overlapping on AS camera. 


Notes:

  • Mode cleaner alignment took some manual tweaking, locked fine around 1k counts. Still no autolocker.
  • At this point, some light was visible on AS and REFL, which was a good sign regarding TTs. 
  • Used green light to align ETMs to support a green 00 mode. 
  • Ensured no recylcying flashes were taking place on AS camera and PRM face camera.
  • Arms were locked using AS55, with the other ITM misalgined, for better SNR than PO[XY]. ASS brought arm powers to ~0.06, which is about what we would expect from 1k MC2 trans instead of 16k.
    • ASS Yarm required debugging, see below.
    • ETMX was getting kicks again. Top Dsub connector on the flange near the ground closer to the end table was a little loose. We should fasten it more securely.
  • At this point, michelson alignment was good. Brought in PRM to see PRC flashes, REFL spot was happy. Brought in SRM to AS sppot. 
  • Saved all optic positions. 
  • Oplevs:
    • PRMs new aligned state is falling off the QPD.
    • ETMs and BS oplev centering are fine, rest are less good, but still on the detector.

 


ASS-RFM issue:

ETMY was not getting its ASC pitch and yaw signals. C1SCY had a red RFM bit (although, it still does now...)

I took a look at the c1rfm simulink diagram and found that C1RFM had an RFM block called C1:RFM-TST_ETMY_[PIT/YAW] and C1SCY had one called C1:TST-SCY_ETMY_[PIT/YAW]. 

It seems that C1TST was illegally being used in a real signal chain, and Jenne's recent work with c1tst broke it. I renamed the channels in C1RFM and C1SCY to C1:RFM-SCY_ETMY_[PIT/YAW], saved, compiled, installed, restarted. All was well.

There are still some  in SCY that have this TST stuff going on, however. They have to do with ALS, it seems, but are SHMEM blocks, not RFM. Namely:

  • C1:TST-SCY_TRY
  • C1:TST-SCY_GLOBALPOS
  • C1:TST-SCY_AMP_CTRL

 

ELOG V3.1.3-